\normalsize



\begin{thebibliography}{20}

\bibitem{cram}
H.~Cram\'er, On the order of magnitude of the difference between consecutive prime numbers. 
{\it Acta Arith.} {\bf 2} (1937), 23--46. 

\bibitem{finch}
S.~R.~Finch, {\it Mathematical Constants}, Cambridge University Press, 2003. 

\bibitem{fischer} 
R.~Fischer, Maximale L\"ucken (Intervallen) von Primzahlenzwillingen, preprint, 2008, 
{\small\tt http://www.fermatquotient.com/PrimLuecken/ZwillingsRekordLuecken}.

\bibitem{fischer2} 
R.~Fischer, Maximale Intervalle von Primzahlenpaaren, web page (in German){. \hfill} \linebreak
Maximal gaps between twin primes are $G_2(p) \approx (1.32032)^{-1}(\log p-(2/3) \log \log p)^3${. \hfill} \linebreak
Maximal gaps between prime triplets are $G_3(p) \approx (2.8582)^{-1}(\log p-(3/4) \log \log p)^4${. \hfill} \linebreak
{\small\tt http://www.fermatquotient.com/PrimLuecken/Max\_Intervalle} (2006).

\bibitem{forbes}
A.~D.~Forbes, Prime $k$-tuplets. Section 21:~List of all possible patterns of prime $k\mbox{-}$tuplets.
The Hardy-Littlewood constants pertaining to the distribution of prime $k\mbox{-}$tuplets, preprint, 2012.  
Available at 
{\small\tt http://anthony.d.forbes.googlepages.com/ktuplets.htm}.

\bibitem{hl}
G.~H.~Hardy and J.~E.~Littlewood, Some problems of `Partitio Numerorum.' III. 
On the expression of a number as a sum of primes. {\it Acta Math.} {\bf 44} (1922), 1--70.

\bibitem{kourbatov2013}
A.~Kourbatov, Maximal gaps between prime $k$-tuples: a statistical approach.
{\it Journal of Integer Sequences}, {\bf 16} (2013), Article 13.5.2. 
{\small\tt http://arxiv.org/abs/1301.2242}

\bibitem{nicely}
T.~R.~Nicely, First occurrence prime gaps, preprint, 2013{. \hfill} \linebreak
Available at {\small\tt http://www.trnicely.net/gaps/gaplist.html}.

\bibitem{toes}
T.\ Oliveira e Silva, S.~Herzog, and S.~Pardi, 
Empirical verification of the even Goldbach conjecture, and computation of prime gaps, up to $4\cdot10^{18}$, 
{\it Math.~Comp.} {\bf 83} %(2014), 
(to appear). 

\bibitem{riesel}
H.~Riesel, {\it Prime Numbers and Computer Methods for Factorization}, Birkh\"auser, 1994. 

\bibitem{rr} 
L. Rodriguez and C. Rivera,
Conjecture 66. Gaps between consecutive twin prime pairs, 2009.
Available at
{\small\tt http://www.primepuzzles.net/conjectures/conj\_066.htm}.

\bibitem{shanks}
D.~Shanks, On maximal gaps between successive primes. {\it Math.~Comp.} {\bf 18} (1964), 646--651.

\bibitem{oeis}
N.~J.~A.~Sloane, {\it The On-Line Encyclopedia of Integer Sequences}, 2012.
Available at {\small\tt http://oeis.org}.

\bibitem{wolf2011} 
M.~Wolf, Some heuristics on the gaps between consecutive primes,
preprint, 2011.  Available at 
{\small\tt http://arxiv.org/abs/1102.0481}.

\bibitem{wolf2001} 
M.~Wolf, Some remarks on the distribution of twin primes,
preprint, 2001.  Available at 
{\small\tt http://arxiv.org/abs/math/0105211}.

\bibitem{zhang}
Y. Zhang, Bounded gaps between primes, {\it Annals of Math.} {\bf 179} %(2014), 
(to appear). 
%
\end{thebibliography}

\documentclass[12pt,reqno]{article}

\usepackage[usenames]{color}
\usepackage{amssymb}
\usepackage{graphicx}
\usepackage{amscd}

\usepackage{color}
\usepackage{fullpage}
\usepackage{float}

\usepackage{graphics,amsmath,amssymb}
\usepackage{amsthm}
\usepackage{amsfonts}
\usepackage{latexsym}
\usepackage{epsf}
\usepackage{changepage}

\begin{document}

\hspace{1mm}

\begin{center}
\vskip 2mm{\LARGE\bf 
Tables of \\
Record Gaps Between Prime Constellations \\
}

\vskip 7mm
\large
Alexei Kourbatov\\
JavaScripter.net/math \\
{\tt akourbatov@gmail.com}
\end{center}

\vskip 6mm

\begin{abstract}
\noindent
We present tables of record gaps between densest prime constellations, or $k$-tuplets.
The tables contain all maximal gaps between prime $k$-tuplets up to $10^{15}$, for each $k\le7$.
\end{abstract}

\vskip 6mm
\noindent
{\bf Introduction.} Mathematicians have long known that there are {\em infinitely many primes}.
We also know that the average gaps between primes near $p$ are about $\log p$ 
(this follows from the {\em prime number theorem}).
However, this knowledge has not helped prove or disprove {\em Cram\'er's conjecture} \cite{cram}
that a record prime gap near $p$ is $O(\log^2{\negthinspace}p)$. 
Computations strongly suggest that $O(\log^2{\negthinspace}p)$ is
a realistic upper bound for the size of prime gaps; moreover, it even seems possible that
the prime gap ending at $p$ is {\bf\em always} less than $\log^2{\negthinspace}p$ \cite[A005250]{nicely,toes,shanks,oeis}. 
Still, the $O(\log^2{\negthinspace}p)$ upper bound for prime gaps remains just a plausible conjecture.

For {\em prime $k$-tuples}, the situation is somewhat similar~--- yet even more uncertain. 
Zhang's theorem \cite{zhang} implies that {\em at least some types of prime $k$-tuples} do occur infinitely often.
The {\em Hardy-Littlewood $k$-tuple conjecture} \cite[p.\,62--68]{hl,riesel} 
suggests that admissible prime $k$-tuples near $x$ occur with average gaps
$a \approx C_k \log^k{\negthinspace}x$; the coefficients $C_k$ 
are reciprocal to the {\em Hardy-Littlewood $k$-tuple constants}\footnote{
For prime $k$-tuplets with $k\le7$, the numerical values are
$C_2\approx0.757392$,
$C_3\approx0.349864$,
$C_4\approx0.240895$,
$C_5\approx0.0986992$,  %4
$C_6\approx0.0578081$,  %1
$C_7\approx0.0185281$.  %4
%For prime $k$-tuplets with $k\ge8$, there may be more than one more constant $C_k$, depending on the specific $k$-tuplet pattern.
%For example, there are three different patterns of prime octuplets corresponding to two different values of $C_8$.
For definitions and further discussion, see also \cite{kourbatov2013}.
} 
\cite{finch,forbes}. 
Consider admissible prime $k$-tuples with a given pattern, and 
denote by $G_k(x)$ the largest gap between $k$-tuples below $x$. 
Then the following heuristic formula expresses 
the expected size of record gaps in terms of the average gap $a$, as $x\to\infty$:
$$
G_k(x) ~\sim~ a \log{x \over a} - ba ~ \lesssim ~  C_k \log^{k+1}{\negthinspace}x, 
 \quad \mbox{ with } \ a = C_k \log^k{\negthinspace}x, \ \ b \approx {2\over k}
 \qquad \mbox{\cite[p.\,5--7]{kourbatov2013}}.
$$
In the absence of rigorous upper bounds for gaps between prime constellations, 
it is important to accumulate extensive numerical evidence on actual record gaps.  
To that end, these tables contain all maximal gaps between prime {\em $k$-tuplets} (the densest admissible clusters of $k$ primes)
up to $10^{15}$, for each $k\le7$. 
Tables for $k=1$, 2, 4 have been previously published by several authors 
\cite{fischer,kourbatov2013,nicely,toes,rr,oeis}.
For $k=1$ the data extends up to $4\times10^{18}$ \cite{toes}.
Tables for $3\le k \le 7$ have been independently obtained as part of my own computation \cite{kourbatov2013,oeis}.
So is it {\bf\em always} true that $G_k(x)<C_k \log^{k+1}{\negthinspace}x$? 
It seems likely; no counterexamples thus far!

\newpage
\section{Record gaps between primes}

\begin{center}TABLE 1 \\Maximal gaps between primes below $4\times10^{18}$ \ 
\cite[OEIS A005250]{nicely,toes,oeis}
\\[0.5em]
\small
\begin{tabular}{rrr|rrr}  
\hline
\multicolumn{2}{r}{Consecutive primes}                            & Gap $g$ & 
\multicolumn{2}{r}{Consecutive primes~\phantom{\fbox{$11111^1$}}} & Gap $g$ \\
[0.5ex]\hline
\vphantom{\fbox{$1^1$}}
          2 &	3	    &   1 &        25056082087 &         25056082543 &  456 \\
          3 &	5	    &   2 &        42652618343 &         42652618807 &  464 \\
          7 &	11	    &   4 &       127976334671 &        127976335139 &  468 \\
         23 &	29	    &   6 &       182226896239 &        182226896713 &  474 \\
         89 &	97	    &   8 &       241160624143 &        241160624629 &  486 \\
        113 &	127	    &  14 &       297501075799 &        297501076289 &  490 \\
        523 &	541	    &  18 &       303371455241 &        303371455741 &  500 \\
        887 &	907	    &  20 &       304599508537 &        304599509051 &  514 \\
       1129 &	1151	    &  22 &       416608695821 &        416608696337 &  516 \\
       1327 &	1361	    &  34 &       461690510011 &        461690510543 &  532 \\
       9551 &	9587	    &  36 &       614487453523 &        614487454057 &  534 \\
      15683 &	15727	    &  44 &       738832927927 &        738832928467 &  540 \\
      19609 &	19661	    &  52 &      1346294310749 &       1346294311331 &  582 \\
      31397 &	31469	    &  72 &      1408695493609 &       1408695494197 &  588 \\
     155921 &	156007	    &  86 &      1968188556461 &       1968188557063 &  602 \\
     360653 &	360749	    &  96 &      2614941710599 &       2614941711251 &  652 \\
     370261 &	370373	    & 112 &      7177162611713 &       7177162612387 &  674 \\
     492113 &	492227	    & 114 &     13829048559701 &      13829048560417 &  716 \\
    1349533 &	1349651	    & 118 &     19581334192423 &      19581334193189 &  766 \\
    1357201 &	1357333	    & 132 &     42842283925351 &      42842283926129 &  778 \\
    2010733 &	2010881	    & 148 &     90874329411493 &      90874329412297 &  804 \\
    4652353 &	4652507	    & 154 &    171231342420521 &     171231342421327 &  806 \\
   17051707 &	17051887    & 180 &    218209405436543 &     218209405437449 &  906 \\
   20831323 &	20831533    & 210 &   1189459969825483 &    1189459969826399 &  916 \\
   47326693 &	47326913    & 220 &   1686994940955803 &    1686994940956727 &  924 \\
  122164747 &	122164969   & 222 &   1693182318746371 &    1693182318747503 & 1132 \\
  189695659 &	189695893   & 234 &  43841547845541059 &   43841547845542243 & 1184 \\
  191912783 &	191913031   & 248 &  55350776431903243 &   55350776431904441 & 1198 \\
  387096133 &	387096383   & 250 &  80873624627234849 &   80873624627236069 & 1220 \\
  436273009 &	436273291   & 282 & 203986478517455989 &  203986478517457213 & 1224 \\
 1294268491 &	1294268779  & 288 & 218034721194214273 &  218034721194215521 & 1248 \\
 1453168141 &	1453168433  & 292 & 305405826521087869 &  305405826521089141 & 1272 \\
 2300942549 &	2300942869  & 320 & 352521223451364323 &  352521223451365651 & 1328 \\
 3842610773 &	3842611109  & 336 & 401429925999153707 &  401429925999155063 & 1356 \\
 4302407359 &	4302407713  & 354 & 418032645936712127 &  418032645936713497 & 1370 \\
10726904659 &	10726905041 & 382 & 804212830686677669 &  804212830686679111 & 1442 \\
20678048297 &	20678048681 & 384 &1425172824437699411 & 1425172824437700887 & 1476 \\
22367084959 &	22367085353 & 394 &                    &                     &      \\
\hline
\end{tabular}
\end{center}
\normalsize

\newpage
\section{Record gaps between twin primes}

\begin{center}TABLE 2 \\Maximal gaps between twin primes $\{p$, $p+2\}$ \ 
\cite[OEIS A113274]{fischer,kourbatov2013,rr,oeis} \\[0.5em]
%\small
\begin{tabular}{rrr|rrr}  
\hline
\multicolumn{2}{r}{Initial primes in twin pairs}                        & Gap $g_2$ & 
\multicolumn{2}{r}{Initial primes in twin pairs~\phantom{\fbox{$1^1$}}} & Gap $g_2$ \\
[0.5ex]\hline
\vphantom{\fbox{$1^1$}}
               3 &               5 &             2 &     24857578817 &     24857585369 &          6552\\
               5 &              11 &             6 &     40253418059 &     40253424707 &          6648\\
              17 &              29 &            12 &     42441715487 &     42441722537 &          7050\\
              41 &              59 &            18 &     43725662621 &     43725670601 &          7980\\
              71 &             101 &            30 &     65095731749 &     65095739789 &          8040\\
             311 &             347 &            36 &    134037421667 &    134037430661 &          8994\\
             347 &             419 &            72 &    198311685749 &    198311695061 &          9312\\
             659 &             809 &           150 &    223093059731 &    223093069049 &          9318\\
            2381 &            2549 &           168 &    353503437239 &    353503447439 &         10200\\
            5879 &            6089 &           210 &    484797803249 &    484797813587 &         10338\\
           13397 &           13679 &           282 &    638432376191 &    638432386859 &         10668\\
           18539 &           18911 &           372 &    784468515221 &    784468525931 &         10710\\
           24419 &           24917 &           498 &    794623899269 &    794623910657 &         11388\\
           62297 &           62927 &           630 &   1246446371789 &   1246446383771 &         11982\\
          187907 &          188831 &           924 &   1344856591289 &   1344856603427 &         12138\\
          687521 &          688451 &           930 &   1496875686461 &   1496875698749 &         12288\\
          688451 &          689459 &          1008 &   2156652267611 &   2156652280241 &         12630\\
          850349 &          851801 &          1452 &   2435613754109 &   2435613767159 &         13050\\
         2868959 &         2870471 &          1512 &   4491437003327 &   4491437017589 &         14262\\
         4869911 &         4871441 &          1530 &  13104143169251 &  13104143183687 &         14436\\
         9923987 &         9925709 &          1722 &  14437327538267 &  14437327553219 &         14952\\
        14656517 &        14658419 &          1902 &  18306891187511 &  18306891202907 &         15396\\
        17382479 &        17384669 &          2190 &  18853633225211 &  18853633240931 &         15720\\
        30752231 &        30754487 &          2256 &  23275487664899 &  23275487681261 &         16362\\
        32822369 &        32825201 &          2832 &  23634280586867 &  23634280603289 &         16422\\
        96894041 &        96896909 &          2868 &  38533601831027 &  38533601847617 &         16590\\
       136283429 &       136286441 &          3012 &  43697538391391 &  43697538408287 &         16896\\
       234966929 &       234970031 &          3102 &  56484333976919 &  56484333994001 &         17082\\
       248641037 &       248644217 &          3180 &  74668675816277 &  74668675834661 &         18384\\
       255949949 &       255953429 &          3480 & 116741875898981 & 116741875918727 &         19746\\
       390817727 &       390821531 &          3804 & 136391104728629 & 136391104748621 &         19992\\
       698542487 &       698547257 &          4770 & 221346439666109 & 221346439686641 &         20532\\
      2466641069 &      2466646361 &          5292 & 353971046703347 & 353971046725277 &         21930\\
      4289385521 &      4289391551 &          6030 & 450811253543219 & 450811253565767 &         22548\\
     19181736269 &     19181742551 &          6282 & 742914612256169 & 742914612279527 &         23358\\
     24215097497 &     24215103971 &          6474 &1121784847637957 & 1121784847661339&         23382\\
\hline
\end{tabular}
\end{center}
\normalsize

\newpage
\section{Record gaps between prime triplets}

\begin{center}TABLE 3.1 \\Maximal gaps between prime triplets $\{p$, $p+2$, $p+6\}$ \ [OEIS A201598] \\[0.5em]
%\small
\begin{tabular}{rrr|rrr}  
\hline
\multicolumn{2}{r}{Initial primes in tuples}                        & Gap $g_3$ & 
\multicolumn{2}{r}{Initial primes in tuples~\phantom{\fbox{$1^1$}}} & Gap $g_3$ \\
[0.5ex]\hline
\vphantom{\fbox{$1^1$}}
        5 &        11 &      6 &      242419361 &       242454281 &  34920 \\
       17 &        41 &     24 &      913183487 &       913222307 &  38820 \\
       41 &       101 &     60 &     1139296721 &      1139336111 &  39390 \\
      107 &       191 &     84 &     2146630637 &      2146672391 &  41754 \\
      347 &       461 &    114 &     2188525331 &      2188568351 &  43020 \\
      461 &       641 &    180 &     3207540881 &      3207585191 &  44310 \\
      881 &      1091 &    210 &     3577586921 &      3577639421 &  52500 \\
     1607 &      1871 &    264 &     7274246711 &      7274318057 &  71346 \\
     2267 &      2657 &    390 &    33115389407 &     33115467521 &  78114 \\
     2687 &      3251 &    564 &    97128744521 &     97128825371 &  80850 \\
     6197 &      6827 &    630 &    99216417017 &     99216500057 &  83040 \\
     6827 &      7877 &   1050 &   103205810327 &    103205893751 &  83424 \\
    39227 &     40427 &   1200 &   133645751381 &    133645853711 & 102330 \\
    46181 &     47711 &   1530 &   373845384527 &    373845494147 & 109620 \\
    56891 &     58907 &   2016 &   412647825677 &    412647937127 & 111450 \\
    83267 &     86111 &   2844 &   413307596957 &    413307728921 & 131964 \\
   167621 &    171047 &   3426 &  1368748574441 &   1368748707197 & 132756 \\
   375251 &    379007 &   3756 &  1862944563707 &   1862944700711 & 137004 \\
   381527 &    385391 &   3864 &  2368150202501 &   2368150349687 & 147186 \\
   549161 &    553097 &   3936 &  2370801522107 &   2370801671081 & 148974 \\
   741677 &    745751 &   4074 &  3710432509181 &   3710432675231 & 166050 \\
   805031 &    809141 &   4110 &  5235737405807 &   5235737580317 & 174510 \\
   931571 &    937661 &   6090 &  8615518909601 &   8615519100521 & 190920 \\
  2095361 &   2103611 &   8250 & 10423696470287 &  10423696665227 & 194940 \\
  2428451 &   2437691 &   9240 & 10660256412977 &  10660256613551 & 200574 \\
  4769111 &   4778381 &   9270 & 11602981439237 &  11602981647011 & 207774 \\
  4938287 &   4948631 &  10344 & 21824373608561 &  21824373830087 & 221526 \\
 12300641 &  12311147 &  10506 & 36385356561077 &  36385356802337 & 241260 \\
 12652457 &  12663191 &  10734 & 81232357111331 &  81232357386611 & 275280 \\
 13430171 &  13441091 &  10920 &186584419495421 & 186584419772321 & 276900 \\
 14094797 &  14107727 &  12930 &297164678680151 & 297164678975621 & 295470 \\
 18074027 &  18089231 &  15204 &428204300934581 & 428204301233081 & 298500 \\
 29480651 &  29500841 &  20190 &450907041535541 & 450907041850547 & 315006 \\
107379731 & 107400017 &  20286 &464151342563471 & 464151342898121 & 334650 \\
138778301 & 138799517 &  21216 &484860391301771 & 484860391645037 & 343266 \\
156377861 & 156403607 &  25746 &666901733009921 & 666901733361947 & 352026 \\
\hline
\end{tabular}
%\medskip
%[OEIS A201598]
%
\end{center}
\normalsize

\newpage
\begin{center}TABLE 3.2  \\Maximal gaps between prime triplets $\{p$, $p+4$, $p+6\}$ \ [OEIS A201596]  % below $10^{15}$
\\[0.5em]
\small
\begin{tabular}{rrr|rrr}  
\hline
\multicolumn{2}{r}{Initial primes in tuples}                        & Gap $g_3$ & 
\multicolumn{2}{r}{Initial primes in tuples~\phantom{\fbox{$1^1$}}} & Gap $g_3$ \\
[0.5ex]\hline
\vphantom{\fbox{$1^1$}}
              7 &         13 &      6 &     2562574867 &      2562620653 &  45786 \\
             13 &         37 &     24 &     2985876133 &      2985923323 &  47190 \\
             37 &         67 &     30 &     4760009587 &      4760057833 &  48246 \\
            103 &        193 &     90 &     5557217797 &      5557277653 &  59856 \\
            307 &        457 &    150 &    10481744677 &     10481806897 &  62220 \\
            457 &        613 &    156 &    19587414277 &     19587476563 &  62286 \\
            613 &        823 &    210 &    25302582667 &     25302648457 &  65790 \\
           2137 &       2377 &    240 &    30944120407 &     30944191387 &  70980 \\
           2377 &       2683 &    306 &    37638900283 &     37638972667 &  72384 \\
           2797 &       3163 &    366 &    49356265723 &     49356340387 &  74664 \\
           3463 &       3847 &    384 &    49428907933 &     49428989167 &  81234 \\
           4783 &       5227 &    444 &    70192637737 &     70192720303 &  82566 \\
           5737 &       6547 &    810 &    74734558567 &     74734648657 &  90090 \\
           9433 &      10267 &    834 &   111228311647 &    111228407113 &  95466 \\
          14557 &      15643 &   1086 &   134100150127 &    134100250717 & 100590 \\
          24103 &      25303 &   1200 &   195126585733 &    195126688957 & 103224 \\
          45817 &      47143 &   1326 &   239527477753 &    239527584553 & 106800 \\
          52177 &      54493 &   2316 &   415890988417 &    415891106857 & 118440 \\
         126487 &     130363 &   3876 &   688823669533 &    688823797237 & 127704 \\
         317587 &     321817 &   4230 &   906056631937 &    906056767327 & 135390 \\
         580687 &     585037 &   4350 &   926175746857 &    926175884923 & 138066 \\
         715873 &     724117 &   8244 &  1157745737047 &   1157745878893 & 141846 \\
        2719663 &    2728543 &   8880 &  1208782895053 &   1208783041927 & 146874 \\
        6227563 &    6237013 &   9450 &  2124064384483 &   2124064533817 & 149334 \\
        8114857 &    8125543 &  10686 &  2543551885573 &   2543552039053 & 153480 \\
       10085623 &   10096573 &  10950 &  4321372168453 &   4321372359523 & 191070 \\
       10137493 &   10149277 &  11784 &  6136808604343 &   6136808803753 & 199410 \\
       18773137 &   18785953 &  12816 & 18292411110217 &  18292411310077 & 199860 \\
       21297553 &   21311107 &  13554 & 19057076066317 &  19057076286553 & 220236 \\
       25291363 &   25306867 &  15504 & 21794613251773 &  21794613477097 & 225324 \\
       43472497 &   43488073 &  15576 & 35806145634613 &  35806145873077 & 238464 \\
       52645423 &   52661677 &  16254 & 75359307977293 &  75359308223467 & 246174 \\
       69718147 &   69734653 &  16506 & 89903831167897 &  89903831419687 & 251790 \\
       80002627 &   80019223 &  16596 &125428917151957 & 125428917432697 & 280740 \\
       89776327 &   89795773 &  19446 &194629563521143 & 194629563808363 & 287220 \\
       90338953 &   90358897 &  19944 &367947033766573 & 367947034079923 & 313350 \\
      109060027 &  109081543 &  21516 &376957618687747 & 376957619020813 & 333066 \\
      148770907 &  148809247 &  38340 &483633763994653 & 483633764339287 & 344634 \\
     1060162843 & 1060202833 &  39990 &539785800105313 & 539785800491887 & 386574 \\
     1327914037 & 1327955593 &  41556 &  \multicolumn{3}{c}{}                     \\
\hline
\end{tabular}

%\medskip
%[OEIS A201596]
\end{center}
\normalsize

\newpage
\section{Record gaps between prime quadruplets}

\begin{center}TABLE 4 \\Maximal gaps between prime quadruplets $\{p$, $p+2$, $p+6$, $p+8\}$ \ [OEIS A113404]
\\[0.5em]
%\small
\begin{tabular}{rrr|rrr}  
\hline
\multicolumn{2}{r}{Initial primes in tuples}                        & Gap $g_4$ & 
\multicolumn{2}{r}{Initial primes in tuples~\phantom{\fbox{$1^1$}}} & Gap $g_4$ \\
[0.5ex]\hline
\vphantom{\fbox{$1^1$}}
         5 &         11 &      6 &      3043111031 &     3043668371 &  557340 \\
        11 &        101 &     90 &      3593321651 &     3593956781 &  635130 \\
       191 &        821 &    630 &      5675642501 &     5676488561 &  846060 \\
       821 &       1481 &    660 &     25346635661 &    25347516191 &  880530 \\
      2081 &       3251 &   1170 &     27329170151 &    27330084401 &  914250 \\
      3461 &       5651 &   2190 &     35643379901 &    35644302761 &  922860 \\
      5651 &       9431 &   3780 &     56390149631 &    56391153821 & 1004190 \\
     25301 &      31721 &   6420 &     60368686121 &    60369756611 & 1070490 \\
     34841 &      43781 &   8940 &     71335575131 &    71336662541 & 1087410 \\
     88811 &      97841 &   9030 &     76427973101 &    76429066451 & 1093350 \\
    122201 &     135461 &  13260 &     87995596391 &    87996794651 & 1198260 \\
    171161 &     187631 &  16470 &     96616771961 &    96618108401 & 1336440 \\
    301991 &     326141 &  24150 &    151023350501 &   151024686971 & 1336470 \\
    739391 &     768191 &  28800 &    164550390671 &   164551739111 & 1348440 \\
   1410971 &    1440581 &  29610 &    171577885181 &   171579255431 & 1370250 \\
   1468631 &    1508621 &  39990 &    210999769991 &   211001269931 & 1499940 \\
   2990831 &    3047411 &  56580 &    260522319641 &   260523870281 & 1550640 \\
   3741161 &    3798071 &  56910 &    342611795411 &   342614346161 & 2550750 \\
   5074871 &    5146481 &  71610 &   1970587668521 &  1970590230311 & 2561790 \\
   5527001 &    5610461 &  83460 &   4231588103921 &  4231591019861 & 2915940 \\
   8926451 &    9020981 &  94530 &   5314235268731 &  5314238192771 & 2924040 \\
  17186591 &   17301041 & 114450 &   7002440794001 &  7002443749661 & 2955660 \\
  21872441 &   22030271 & 157830 &   8547351574961 &  8547354997451 & 3422490 \\
  47615831 &   47774891 & 159060 &  15114108020021 & 15114111476741 & 3456720 \\
  66714671 &   66885851 & 171180 &  16837633318811 & 16837637203481 & 3884670 \\
  76384661 &   76562021 & 177360 &  30709975578251 & 30709979806601 & 4228350 \\
  87607361 &   87797861 & 190500 &  43785651890171 & 43785656428091 & 4537920 \\
 122033201 &  122231111 & 197910 &  47998980412211 & 47998985015621 & 4603410 \\
 132574061 &  132842111 & 268050 &  55341128536691 & 55341133421591 & 4884900 \\
 204335771 &  204651611 & 315840 &  92944027480721 & 92944033332041 & 5851320 \\
 628246181 &  628641701 & 395520 & 412724560672211 &412724567171921 & 6499710 \\
1749443741 & 1749878981 & 435240 & 473020890377921 &473020896922661 & 6544740 \\
2115383651 & 2115824561 & 440910 & 885441677887301 &885441684455891 & 6568590 \\
2128346411 & 2128859981 & 513570 & 947465687782631 &947465694532961 & 6750330 \\
2625166541 & 2625702551 & 536010 & 979876637827721 &979876644811451 & 6983730 \\
2932936421 & 2933475731 & 539310 &                 &                &         \\
\hline
\end{tabular}

%\medskip
%[OEIS A113404]
%
\end{center}
\normalsize

\newpage
\section{Record gaps between prime quintuplets}

\begin{center}TABLE 5.1 \\
Maximal gaps between prime quintuplets $\{p$, $p+2$, $p+6$, $p+8$, $p+12\}$  \\[0.5em]
%\small
\begin{tabular}{rrr|rrr}  
\hline
\multicolumn{2}{r}{Initial primes in tuples}                        & Gap $g_5$ & 
\multicolumn{2}{r}{Initial primes in tuples~\phantom{\fbox{$1^1$}}} & Gap $g_5$ \\
[0.5ex]\hline
\vphantom{\fbox{$1^1$}}
              5 &             11 &           6 &    107604759671 &   107616100511 &    11340840 \\
             11 &            101 &          90 &    140760439991 &   140772689501 &    12249510 \\
            101 &           1481 &        1380 &    162661360481 &   162673773671 &    12413190 \\
           1481 &          16061 &       14580 &    187735329491 &   187749510491 &    14181000 \\
          22271 &          43781 &       21510 &    327978626531 &   327994719461 &    16092930 \\
          55331 &         144161 &       88830 &    508259311991 &   508275672341 &    16360350 \\
         536441 &         633461 &       97020 &    620537349191 &   620554105931 &    16756740 \\
         661091 &         768191 &      107100 &    667672901711 &   667689883031 &    16981320 \\
        1461401 &        1573541 &      112140 &   1079628551621 &  1079646141851 &    17590230 \\
        1615841 &        1917731 &      301890 &   1104604933841 &  1104624218981 &    19285140 \\
        5527001 &        5928821 &      401820 &   1182148717481 &  1182168243071 &    19525590 \\
       11086841 &       11664551 &      577710 &   1197151034531 &  1197173264711 &    22230180 \\
       35240321 &       35930171 &      689850 &   2286697462781 &  2286720012251 &    22549470 \\
       53266391 &       54112601 &      846210 &   2435950632251 &  2435980618781 &    29986530 \\
       72610121 &       73467131 &      857010 &   3276773115431 &  3276805283951 &    32168520 \\
       92202821 &       93188981 &      986160 &   5229301162991 &  5229337555061 &    36392070 \\
      117458981 &      119114111 &     1655130 &   9196865051651 &  9196903746881 &    38695230 \\
      196091171 &      198126911 &     2035740 &  14660925945221 & 14660966101421 &    40156200 \\
      636118781 &      638385101 &     2266320 &  21006417451961 & 21006458070461 &    40618500 \\
      975348161 &      977815451 &     2467290 &  22175175736991 & 22175216733491 &    40996500 \\
     1156096301 &     1158711011 &     2614710 &  22726966063091 & 22727007515411 &    41452320 \\
     1277816921 &     1281122231 &     3305310 &  22931291089451 & 22931338667591 &    47578140 \\
     1347962381 &     1351492601 &     3530220 &  31060723328351 & 31060771959221 &    48630870 \\
     2195593481 &     2199473531 &     3880050 &  85489258071311 & 85489313115881 &    55044570 \\
     3128295551 &     3132180971 &     3885420 &  90913430825291 & 90913489290971 &    58465680 \\
     4015046591 &     4020337031 &     5290440 &  96730325054171 & 96730390102391 &    65048220 \\
     8280668651 &     8286382451 &     5713800 & 199672700175071 &199672765913051 &    65737980 \\
     9027127091 &     9033176981 &     6049890 & 275444947505591 &275445018294491 &    70788900 \\
    15686967971 &    15693096311 &     6128340 & 331992774272981 &331992848243801 &    73970820 \\
    18901038971 &    18908988011 &     7949040 & 465968834865971 &465968914851101 &    79985130 \\
    21785624291 &    21793595561 &     7971270 & 686535413263871 &686535495684161 &    82420290 \\
    30310287431 &    30321057581 &    10770150 & 761914822198961 &761914910291531 &    88092570 \\
\hline
\end{tabular}

\medskip
[OEIS A201073]
\end{center}
\normalsize

\newpage
\begin{center}TABLE 5.2 \\Maximal gaps between prime quintuplets $\{p$, $p+4$, $p+6$, $p+10$, $p+12\}$ \\[0.5em]
%\small
\begin{tabular}{rrr|rrr}  
\hline
\multicolumn{2}{r}{Initial primes in tuples}                        & Gap $g_5$ & 
\multicolumn{2}{r}{Initial primes in tuples~\phantom{\fbox{$1^1$}}} & Gap $g_5$ \\
[0.5ex]\hline
\vphantom{\fbox{$1^1$}}
              7 &             97 &          90 &     15434185927 &     15440743597 &     6557670\\
             97 &           1867 &        1770 &     17375054227 &     17381644867 &     6590640\\
           3457 &           5647 &        2190 &     17537596327 &     17544955777 &     7359450\\
           5647 &          15727 &       10080 &     25988605537 &     25997279377 &     8673840\\
          19417 &          43777 &       24360 &     66407160637 &     66416495137 &     9334500\\
          43777 &          79687 &       35910 &     74862035617 &     74871605947 &     9570330\\
         101107 &         257857 &      156750 &     77710388047 &     77723371717 &    12983670\\
        1621717 &        1830337 &      208620 &    144124106167 &    144138703987 &    14597820\\
        3690517 &        3995437 &      304920 &    210222262087 &    210238658797 &    16396710\\
        5425747 &        5732137 &      306390 &    585234882097 &    585252521167 &    17639070\\
        8799607 &        9127627 &      328020 &    926017532047 &    926036335117 &    18803070\\
        9511417 &        9933607 &      422190 &    986089952917 &    986113345747 &    23392830\\
       16388917 &       16915267 &      526350 &   2819808136417 &   2819832258697 &    24122280\\
       22678417 &       23317747 &      639330 &   3013422626107 &   3013449379477 &    26753370\\
       31875577 &       32582437 &      706860 &   3538026326827 &   3538053196957 &    26870130\\
       37162117 &       38028577 &      866460 &   4674635167747 &   4674662545867 &    27378120\\
       64210117 &       65240887 &     1030770 &   5757142722757 &   5757171559957 &    28837200\\
      119732017 &      120843637 &     1111620 &   7464931087717 &   7464961813867 &    30726150\\
      200271517 &      201418957 &     1147440 &   8402871269197 &   8402904566467 &    33297270\\
      203169007 &      204320107 &     1151100 &   9292699799017 &   9292733288557 &    33489540\\
      241307107 &      242754637 &     1447530 &  10985205390997 &  10985239010737 &    33619740\\
      342235627 &      344005297 &     1769670 &  12992848206847 &  12992884792957 &    36586110\\
      367358347 &      369151417 &     1793070 &  15589051692667 &  15589094176627 &    42483960\\
      378200227 &      380224837 &     2024610 &  24096376903597 &  24096421071127 &    44167530\\
      438140947 &      440461117 &     2320170 &  37371241083097 &  37371285854467 &    44771370\\
      446609407 &      448944487 &     2335080 &  38728669335607 &  38728728308527 &    58972920\\
      711616897 &      714020467 &     2403570 &  91572717670537 &  91572784840627 &    67170090\\
      966813007 &      970371037 &     3558030 & 109950817237357 & 109950886775827 &    69538470\\
     2044014607 &     2048210107 &     4195500 & 325554440818297 & 325554513360487 &    72542190\\
     3510456787 &     3514919917 &     4463130 & 481567288596127 & 481567361629087 &    73032960\\
     4700738167 &     4705340527 &     4602360 & 501796510663237 & 501796584764467 &    74101230\\
     5798359657 &     5803569847 &     5210190 & 535243109721577 & 535243185965557 &    76243980\\
     7896734467 &     7902065527 &     5331060 & 657351798174427 & 657351876771637 &    78597210\\
    12654304207 &    12659672737 &     5368530 & 818872754682547 & 818872840949077 &    86266530\\
    13890542377 &    13896088897 &     5546520 & 991851356676277 & 991851464273767 &   107597490\\
    14662830817 &    14668797037 &     5966220 &                 &                 &            \\
\hline
\end{tabular}

\medskip
[OEIS A201062]
\end{center}
\normalsize

\newpage
\section{Record gaps between prime sextuplets}
\begin{adjustwidth}{-9mm}{}
\begin{center}TABLE 6 \\Maximal gaps between prime sextuplets $\{p$, $p+4$, $p+6$, $p+10$, $p+12$, $p+16\}$ \\[0.5em]
\normalsize
\begin{tabular}{rrr|rrr}  
\hline
\multicolumn{2}{r}{Initial primes in tuples}                        & Gap $g_6$ & 
\multicolumn{2}{r}{Initial primes in tuples~\phantom{\fbox{$1^1$}}} & Gap $g_6$ \\
[0.5ex]\hline
\vphantom{\fbox{$1^1$}}
               7 &            97 &            90 &    422088931207 &   422248594837 &     159663630\\
              97 &         16057 &         15960 &    427190088877 &   427372467157 &     182378280\\
           19417 &         43777 &         24360 &    610418426197 &   610613084437 &     194658240\\
           43777 &       1091257 &       1047480 &    659829553837 &   660044815597 &     215261760\\
         3400207 &       6005887 &       2605680 &    660863670277 &   661094353807 &     230683530\\
        11664547 &      14520547 &       2856000 &    853633486957 &   853878823867 &     245336910\\
        37055647 &      40660717 &       3605070 &   1089611097007 &  1089869218717 &     258121710\\
        82984537 &      87423097 &       4438560 &   1247852774797 &  1248116512537 &     263737740\\
        89483827 &      94752727 &       5268900 &   1475007144967 &  1475318162947 &     311017980\\
        94752727 &     112710877 &      17958150 &   1914335271127 &  1914657823357 &     322552230\\
       381674467 &     403629757 &      21955290 &   1953892356667 &  1954234803877 &     342447210\\
      1569747997 &    1593658597 &      23910600 &   3428196061177 &  3428617938787 &     421877610\\
      2019957337 &    2057241997 &      37284660 &   9367921374937 &  9368397372277 &     475997340\\
      5892947647 &    5933145847 &      40198200 &  10254799647007 & 10255307592697 &     507945690\\
      6797589427 &    6860027887 &      62438460 &  13786576306957 & 13787085608827 &     509301870\\
     14048370097 &   14112464617 &      64094520 &  21016714812547 & 21017344353277 &     629540730\\
     23438578897 &   23504713147 &      66134250 &  33157788914347 & 33158448531067 &     659616720\\
     24649559647 &   24720149677 &      70590030 &  41348577354307 & 41349374379487 &     797025180\\
     29637700987 &   29715350377 &      77649390 &  72702520226377 & 72703333384387 &     813158010\\
     29869155847 &   29952516817 &      83360970 &  89165783669857 & 89166606828697 &     823158840\\
     45555183127 &   45645253597 &      90070470 & 122421000846367 &122421855415957 &     854569590\\
     52993564567 &   53086708387 &      93143820 & 139864197232927 &139865086163977 &     888931050\\
     58430706067 &   58528934197 &      98228130 & 147693859139077 &147694869231727 &    1010092650\\
     93378527647 &   93495691687 &     117164040 & 186009633998047 &186010652137897 &    1018139850\\
     97236244657 &   97367556817 &     131312160 & 202607131405027 &202608270995227 &    1139590200\\
    240065351077 &  240216429907 &     151078830 & 332396845335547 &332397997564807 &    1152229260\\
    413974098817 &  414129003637 &     154904820 & 424681656944257 &424682861904937 &    1204960680\\
    419322931117 &  419481585697 &     158654580 & 437804272277497 &437805730243237 &    1457965740\\
\hline
\end{tabular}

\medskip
[OEIS A200503]
\end{center}
\end{adjustwidth}
\normalsize

\phantom{\vspace{1cm}}

\newpage
\section{Record gaps between prime septuplets}

\begin{adjustwidth}{-9mm}{}
\begin{center}TABLE 7.1 
\\Maximal gaps between prime 7-tuples $\{p$, $p+2$, $p+8$, $p+12$, $p+14$, $p+18$, $p+20\}$ \\[0.5em]
%\small
\begin{tabular}{rrr|rrr}  
\hline
\multicolumn{2}{r}{Initial primes in tuples}                        & Gap $g_7$ & 
\multicolumn{2}{r}{Initial primes in tuples~\phantom{\fbox{$1^1$}}} & Gap $g_7$ \\
[0.5ex]\hline
\vphantom{\fbox{$1^1$}}
           5639 &          88799 &       83160 &   1554893017199 &  1556874482069 &  1981464870\\
          88799 &         284729 &      195930 &   2088869793539 &  2090982626639 &  2112833100\\
         284729 &         626609 &      341880 &   2104286376329 &  2106411289049 &  2124912720\\
        1146779 &        6560999 &     5414220 &   2704298257469 &  2706823007879 &  2524750410\\
        8573429 &       17843459 &     9270030 &   3550904257709 &  3553467600029 &  2563342320\\
       24001709 &       42981929 &    18980220 &   4438966968419 &  4442670730019 &  3703761600\\
       43534019 &       69156539 &    25622520 &   9996858589169 & 10000866474869 &  4007885700\\
       87988709 &      124066079 &    36077370 &  21937527068909 & 21942038052029 &  4510983120\\
      157131419 &      208729049 &    51597630 &  29984058230039 & 29988742571309 &  4684341270\\
      522911099 &      615095849 &    92184750 &  30136375346249 & 30141383681399 &  5008335150\\
      706620359 &      832143449 &   125523090 &  32779504324739 & 32784963061379 &  5458736640\\
     1590008669 &     1730416139 &   140407470 &  40372176609629 & 40377635870639 &  5459261010\\
     2346221399 &     2488117769 &   141896370 &  42762127106969 & 42767665407989 &  5538301020\\
     3357195209 &     3693221669 &   336026460 &  54620176867169 & 54626029928999 &  5853061830\\
    11768282159 &    12171651629 &   403369470 &  63358011407219 & 63365153990639 &  7142583420\\
    30717348029 &    31152738299 &   435390270 &  79763188368959 & 79770583970249 &  7395601290\\
    33788417009 &    34230869579 &   442452570 & 109974651670769 &109982176374599 &  7524703830\\
    62923039169 &    63550891499 &   627852330 & 145568747217989 &145576919193689 &  8171975700\\
    68673910169 &    69428293379 &   754383210 & 196317277557209 &196325706400709 &  8428843500\\
    88850237459 &    89858819579 &  1008582120 & 221953318490999 &221961886287509 &  8567796510\\
   163288980299 &   164310445289 &  1021464990 & 249376874266769 &249385995968099 &  9121701330\\
   196782371699 &   197856064319 &  1073692620 & 290608782523049 &290618408585369 &  9626062320\\
   421204876439 &   422293025249 &  1088148810 & 310213774327979 &310225023265889 & 11248937910\\
   427478111309 &   428623448159 &  1145336850 & 471088826892779 &471100312066829 & 11485174050\\
   487635377219 &   489203880029 &  1568502810 & 631565753063879 &631578724265759 & 12971201880\\
   994838839439 &   996670266659 &  1831427220 & 665514714418439 &665530090367279 & 15375948840\\
\hline
\end{tabular}

\medskip
[OEIS A201251]
\end{center}
\end{adjustwidth}
\normalsize

\newpage
\begin{adjustwidth}{-9mm}{}
\begin{center}TABLE 7.2 \\
Maximal gaps between prime 7-tuples $\{p$, $p+2$, $p+6$, $p+8$, $p+12$, $p+18$, $p+20\}$  \\[0.5em]
%\small
\begin{tabular}{rrr|rrr}  
\hline
\multicolumn{2}{r}{Initial primes in tuples}                        & Gap $g_7$ & 
\multicolumn{2}{r}{Initial primes in tuples~\phantom{\fbox{$1^1$}}} & Gap $g_7$ \\
[0.5ex]\hline
\vphantom{\fbox{$1^1$}}
             11 &         165701 &     165690 &   382631592641 &   383960791211 & 1329198570\\
         165701 &        1068701 &     903000 &   711854781551 &   714031248641 & 2176467090\\
        1068701 &       11900501 &   10831800 &  2879574595811 &  2881987944371 & 2413348560\\
       25658441 &       39431921 &   13773480 &  3379186846151 &  3381911721101 & 2724874950\\
       45002591 &       67816361 &   22813770 &  5102247756491 &  5105053487531 & 2805731040\\
       93625991 &      124716071 &   31090080 &  5987254671311 &  5990491102691 & 3236431380\\
      257016491 &      300768311 &   43751820 &  7853481899561 &  7857040317011 & 3558417450\\
      367438061 &      428319371 &   60881310 & 11824063534091 & 11828142800471 & 4079266380\\
      575226131 &      661972301 &   86746170 & 16348094430581 & 16353374758991 & 5280328410\\
     1228244651 &     1346761511 &  118516860 & 44226969237161 & 44233058406611 & 6089169450\\
     1459270271 &     1699221521 &  239951250 & 54763336591961 & 54771443197181 & 8106605220\\
     2923666841 &     3205239881 &  281573040 &154325181803321 &154333374270191 & 8192466870\\
    10180589591 &    10540522241 &  359932650 &157436722520921 &157445120715341 & 8398194420\\
    15821203241 &    16206106991 &  384903750 &281057032201481 &281065611322031 & 8579120550\\
    23393094071 &    23911479071 &  518385000 &294887168565161 &294896169845351 & 9001280190\\
    37846533071 &    38749334621 &  902801550 &309902902299701 &309914040972071 &11138672370\\
   158303571521 &   159330579041 & 1027007520 &419341934631071 &419354153220461 &12218589390\\
   350060308511 &   351146640191 & 1086331680 &854077393259801 &854090557643621 &13164383820\\
\hline
\end{tabular}

\medskip
[OEIS A201051]
\end{center}
\end{adjustwidth}

\bigskip
\hrule
\bigskip

\noindent 2010 {\it Mathematics Subject Classification}: 11N05
%Primary 11N05; Secondary 60G70.

\noindent \emph{Keywords: } 
distribution of primes, prime $k$-tuple, Hardy-Littlewood conjecture,
extreme value statistics, prime gap, Cram\'er conjecture, 
prime constellation, twin prime conjecture, prime triplet, prime quadruplet, 
prime quintuplet, prime sextuplet, prime septuplet.

\bigskip
\hrule
\bigskip

\noindent (Concerned with OEIS sequences
A005250,
A113274,
A113404,
A200503,
A201051,
A201062,
A201073,
A201251,
A201596,
A201598,
A202281,
A202361.)

\bigskip

\end{document}